OpenStudy (anonymous):

simplify . 13[6^2 / (5^2 - 4^2) + 9 ]

OpenStudy (anonymous):

\[13(\frac{6^2}{(5^2-4^2)+9})\] \[13(\frac{36}{(25-16)+9})\] \[13(\frac{36}{9+9})\] \[13(\frac{36}{18})\] \[13(2)\] \[26\]